OpenStudy (anonymous):

how do you find the direction of u in the direction derivative function if it's not given?

OpenStudy (anonymous):

The problem states Consider the function f=2x^2+xy at Po(3,-1) I got Fx=4x+4 and Fy=2+x but i don't know what <a,b,c> is